together danny and luis have 40 marbles, danny has 8 more than luis, how many marbles does danny have?
Answer by stanbon(75887)

together danny and luis have 40 marbles, danny has 8 more than luis, how many marbles does danny have?
------
Equation:
D + L = 40 marbles
D = L + 8
----------
Substitute for "D" and solve for "L"::
L + 8 + L = 40
2L = 32
L = 16 (# Luis has)
---
Solve for "D":
D = L + 8
D = 16 + 8 = 24 (# Danny has)
===============================
Cheers,
Stan H.
======================